In our population the incidence of polyposis of the large intestine is 1:8,000. The disorder is determined by the autosomal dominant allele. What is the frequency of disease-causing allele in the population?
1
Expert's answer
2020-04-08T13:34:58-0400

According to Hardy–Weinberg principle p2 + 2pq + q2 = 100% = 1,

where, p2 is the frequency of the AA genotype,

2pq - frequency of the genotype Aa,

q2 is the frequency of the aa genotype.

Then q2 = 1 - (p2 + 2pq)

q2 = 1 - (1:8,000) = 0,99

q = 0,99

p = 1 - 0,99 = 0,01

Frequency of disease-causing allele in the population is 0,01.